OpenStudy (anonymous):

9. Which function family would be used to solve the following question? Chuck and his friends are taking a road trip. If they have traveled 224 miles in 4 hours, what is their average speed? Exponential Function Family Trigonometric Function Family Linear Function Family Logarithmic Function Famil

OpenStudy (anonymous):

Linear

OpenStudy (anonymous):

The average speed is: AvSpeed = (Total distance)/(Time consumed) = 224/4 = 56 miles/hour You use only division, so it's linear

OpenStudy (anonymous):

well, it is Speed = distance / time. 224/4 = 56 therefore, it being linear

OpenStudy (anonymous):

thanks btw your one year older than!